One of two positive integers is 8 more than the other. If the product of the two integers is 84, find the integers. Must be solved in an algebraic solution.
Found 2 solutions by richwmiller, edjones:
Answer by richwmiller(17219)

let n= first and n+8= second
n*(n+8)=84
n^2+8n=84
complete the square
8/2=4
4^2=16
add 16 to both sides
n^2+8n+16=84+16
(n+4)^2=100
n+4=+\-10
n=-4+10
n=-4-10
n=6
n=-14
we discard the negative root
n=6
n+8=14
check
6*14=84
84=ok

Answer by edjones(8007)

Let the numbers be x and x+8
x(x+8)=84
x^2+8x-84=0
(x+14)(x-6)=0
x=6, x+8=14
.
Ed